Patrick is five times as old as ben.after four years patrick will be three times as old as ben.what are their present ages?
Answer by htmentor(1343)

Let B = Ben's age now
Then Patricks's age now = 5B
After 4 years, Ben will be B+4 and Patrick will be 5B+4, and 3 times as old as Ben:
5B+4 = 3(B+4)
Solve for B:
5B + 4 = 3B + 12
2B = 8
B = 4
So Ben is 4 yrs old and Patrick is 20 yrs old